OpenStudy (anonymous):

5 - 2 10-2x ---- = ------ x-6 x-6 I want to see if I got the same answer

OpenStudy (anonymous):

The 5 at the beginning of the equation is a whole number by itself .only the 2 over the x-6 is a fraction

jhonyy9 (jhonyy9):

2 10 - 2x 5 - ----- = ------- --- 5(x-6) - 2 = 10 - 2x x-6 x-6 5x -30 -2 =10 - 2x 7x = 42 x=42/7 =6 x=6

jhonyy9 (jhonyy9):

but how you can seeing x dont possible being 6 because than the denominator will be zero

OpenStudy (anonymous):

x is not equal to 6 so ther is no solution set right?

jhonyy9 (jhonyy9):

check it please newly because if i have wrote correct your equation with this fractions the result is x=6

OpenStudy (anonymous):

okay! thanks oops i know I confused you that is the next part of the answer sheet I have. Sorry!!!

jhonyy9 (jhonyy9):

ok